D/E Night UT-07-Feb-2020

Plan

OSA: J. Williams
Observers: S. Walsh, J. Hill, D. Miller, A. Conrad
ISA: O. Kuhn

  • Start of night PEPSI faint object mode testing - 2hours (handover to LBTO staff at 9:20 PM local time)
  • Geo-Sync satellite - Anik F1R - requires IRTC on SX w/SOUL AO - pending release of IRTC (and computer issues solved)
  • TCS + MCSPU 2020B release - testing occurs during regular use of telescope - new patch - test during night
  • LUCI Photometric standards only if clear
  • Mixed-mode LBC/LUCI tests - dithering tests - medium priority - 0.5 hr - presets worked in full binocular mode
  • Some follow up on ITs:
    • Make sure LBCs can really guide. We know BTech is running and tech chip images were taken closed dome. Don't expect any surprises open dome, but perhaps we should test?
    • DONE already: Make sure offsets made via telsvc gui are noted as completed by telsvc gui (there was an IT that this the offset got stuck a while back, but another note/email saying it was working fine). Would not hurt to check.
  • D-time targets
    • Cals needed for LUCI-2 G210 K-band 0.5" slit CWL 2.27 microns (Rothberg poor weather program)
      • Location: RM507-5/4: /home/observer/share/Service_Observing/Eng_Jan_2020/Rothberg_Local_Mergers/LUCI_CLosed_Dome_Cals/Cals_G210_0pt5_CWL2pt27_L2.xml
    • Quasar Q1237+146 with MODS 1 hr before twilight (with Moon) and with LUCI @ G200 zJ before.

Log

All times in MST

Opened and started with binocular PEPSI.

20:00 Josh has been struggling with pointing and collimating since we opened. He has been on the phone with John H for a while now.

20:00 LBCR Trackers will not come up. There is a problem with the voltage on line 3: "bad input voltage line 3 (+19.95V instead of +18.5V+-1.0V)" Critical IT 8048 opened. I tried to turn on the LBC systems three times, always with the same error. I'll continue without the trackers for now.

20:00 MODS2 test images taken again in dual, red-only and blue-only. Spot is on MODS2 dichroic and not on field lens and has been there since September. See IT 7798.

20:25 Test biases with both LBCs. Blue Chip 2 is noisy. Temp = -2.0 C.

20:30 Aligned LUCI2 with N30 + N30_FS. Took a long time to get stretch right on Aladin to see spots (Ks, DIT=10s, NDIT=3-10) and, even then, it did not seem possible to align all 4 corners. Good enough and now I'm moving the mask out.

20:45 DX AdSec has been going into Failure mode several times already. Shane and Doug are looking into it.

20:53 Josh can get a star on one side, then when he saves IE/CA values and when he spirals to find it on the other side and then restores the IE/CA values on the first side, the star does not come back. Not much progress is being made and the suggestion was to switch to MODS or any instrument that doesn't use the tertiary. (We could do LBC + MODS tests, but there's not much to be done with MODS observing in close-to-full moon).

21:00 Switching to 2020B.

21:40 Switching to LUCI2 and IRTC. Doug, John, Al and Shane are here. Pointed worked for bino LUCI (IRTC + LUCI2). Odd... Was this due to thermal effects on an already difficult to copoint bino PEPSIPFU? Or something about the switch of TCS build (not the build itself, but restarting of a subsystem)? Would be worth trying to authorize bino PEPSIPFU once again at the end of the night.

22:30 Doug and Shane are working on DX AdSec; it doesn't want to stay set.

22:30 Tested offsets via the telsvc GUI --- do they work in the patched 2020B? Yes. This also further tested the problem reported in IT 7965. telsvc GUI is no longer getting stuck on "Commit".

22:30 Josh is sending some active presets to off-axis stars to exercise the new build a little more.

22:35 Bright spot in middle of SX pupil... it just appeared on left_wfscsubimage000049.fits and did not seem to be there earlier, but earlier stars may have been brighter.

22:50 Shane found that the ARGOS light source was "probably ON" and the shutter was open. Shane turned it "probably" OFF and closed the shutter.

22:55 Doug reverting to UAO to see if there are still problems with the keeping the DX AdSec set.

01:15 Since about 23h, Doug and Shane have sent presets using IDL scripts, closed the loop on SX and taken IRTC images. But the DX AdSec goes into failure mode and was left rested (while keeping an eye on the wind speed). Doug is on the phone with Guido. This is similar to IT 7976. At the end of the night, Doug submitted critical IT 8051.

01:25 SX M1 panicked.

02:15 Sending a binocular preset: AO on SX and track on DX; to ANIK. Used paddle to move object around on LUCI. With object at 1023, 1110.5 on LUCI2 (luci2.20200207.0017.fits), it is in IRTC. But, there are two similar objects that are separated N-S by about 28.8". The first image after the slew was luci2.20200207.0014.fits the object is at 853, 1162. Taking a few more luci2 images at BrG and then PaB as the IRTC images come through at H and J. The three objects on the LUCI FOV move relative to one another, but they are (well, not round on LUCI2 because the DX shell is rested) and on frame 17 there is a trailed star.

03:25 Going to a PSF star: V=6.6. Anik has V=8 - color V-H = 1.4 (like G2 star). It appears on the images to be brighter than the satellite.

04:25 Going to another of the ANIK satellites. The one that was to the N of the one we observed earlier (the northernmost of the 3 in image 0014). Now, in the initial acq image 0028, there are 3 objects along a diagonal from top left to bottom right. We paddled the one at the top left, the brightest one, to the "hotspot" where it was seen by IRTC. The final LUCI2 acq image is 0043 and the object is at 992, 1105 on it.

05:30 Going to a PSF star for this second satellite. Natural seeing is getting worse now. Problems getting this second star. Giving up and going to reconfig for PEPSI PFU.

05:55 Reconfig for PEPSIPFU binocular to try to acquire stars on both guiders, now that the telescope is in better thermal equilibrium (at least one potential factor in the problems earlier tonight). Josh says that when he authorized the PEPSI PFUs now, he saw the same following errors as he did yesterday morning, but this time he was able to clear them. He tried a preset, but realized that there are probably some things in the beam after Ilya shut down PEPSI earlier tonight and we won't be able to make this test now.

06:30 Closing up.

06:35 Running Cals_G210_Opt5_CWL2pt27_L2.xml *luci2.20200207.0052-56 lamp off flats
  • 0057-61 lamp on flats
  • 0062-0063 60-sec darks for arcs
  • 0064-0065 60-sec Argon
  • 0066-0067 60-sec Xenon
  • 0068-0069 60-sec Krypton

-- BarryRothberg - 06 Feb 2020
Topic revision: r8 - 07 Feb 2020, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback